Solve the following

π‘Ž) Maximize 𝑓(π‘₯, 𝑦) = 2π‘₯ + 5𝑦; subject to the constraints.

2𝑦 βˆ’ π‘₯ ≀ 8; π‘₯ βˆ’ 𝑦 ≀ 4; π‘₯ β‰₯ 0; 𝑦 β‰₯ 0.

𝑏) Minimize 𝑧 = 3π‘₯ + 𝑦; subject to the constraints.

3π‘₯ + 5𝑦 β‰₯ 15; π‘₯ + 6𝑦 β‰₯ 9; π‘₯ β‰₯ 0; 𝑦 β‰₯ 0.


1
Expert's answer
2021-05-07T09:32:34-0400

a)

"f(x, y)=2x+5y"

"OA: x=0, 0\leq y\leq4"


"f(0, y)=5y"

"f(0, 0)=0, f(0, 4)=20"


"AB: 2y-x=8=>x=2y-8, 4\leq y\leq12"

"f(2y-8, y)=2(2y-8)+5y=9y-16"

"f(0, 4)=20, f(16, 12)=92"



"BC: x-y=4=>x=y+4, 0\leq y\leq12"

"f(y+4, y)=2(y+4)+5y=7y+8"

"f(16, 12)=92, f(4, 0)=8"



"CO: y=0, 0\leq x\leq4"

"f(x, 0)=2x"

"f(0, 0)=0, f(4,0)=8"



The function "f(x, y)" has maximum with value of "92" at "x=16, y=12"

subject to the constraints.


b)




"z=3x+y"


"\begin{matrix}
 3x+5y=15 \\
 x+6y=9
\end{matrix}"

"x=\dfrac{45}{13}, y=\dfrac{12}{13}"


"x=0, y\geq3"


"z=y\geq3"

"z(0, 3)=3"

"AB: 3x+5y=15=>x=5-\dfrac{5}{3}y, \dfrac{12}{13}\leq y\leq3"

"z=15-5y+y=15-4y"

"z(0, 3)=3, z(\dfrac{45}{13}, \dfrac{12}{13})=\dfrac{147}{13}"



"BC: x+6y=9=>x=9-6y, 0\leq y\leq\dfrac{12}{13}"

"z=27-18y+y=27-17y"

"z(\dfrac{45}{13}, \dfrac{12}{13})=\dfrac{147}{13}, z(9, 0)=27"



"y=0, x\geq9"


"z=3x\geq27"

The function "z" has minimum with value of "3" at "x=0, y=3"

subject to the constraints.